# Ask, and it shall be given: Turing completeness of prompting

### v1

@@ 1 Introduction
Nevertheless, existing theoretical studies fail to align with the LLM prompting practice, i.e., the one-model-many-tasks paradigm. In the era of LLMs, people train a single general-purpose LLM and provide the LLM with different prompts to perform different tasks. Since the success of GPT (Brown et al., 2020), the LLM prompting paradigm has revolutionized machine learning (Liu et al., 2023b). For example, a bonus capability arising from prompting is zero-shot learning (Wei et al., 2022a): when provided with suitable prompts, LLMs can even perform novel tasks that are not present in their training corpora. Such empirical success calls for a theoretical understanding of the LLM prompting paradigm:
@@
We answer this call and present the first theory on the LLM prompting paradigm to the best of our knowledge. In this work, we show that prompting is in fact Turing-complete: there exists a finite-size Transformer such that for any computable function, there exists a corresponding prompt following which the Transformer computes the function. Furthermore, we show that prompting is not only universal but also efficiently universal: even though we only use a single finite-size Transformer, it can still achieve nearly the same complexity bounds as that of the class of all unbounded-size Transformers.
@@
  • CoT complexity. Our $\varGamma$ can compute any $\mathsf{TIME}_{2}(t(n))$ function within $\mathrm{O}(t(n))$ CoT steps and can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(t(n)\log t(n))$ CoT steps for any length-$n$ input. Notably, our result shows that even a single Transformer can still achieve nearly the same CoT complexity as the class of all Transformers does.
  • Precision complexity. Our constructed $\varGamma$ can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(\log(n+t(n)))$ bits of precision for any length-$n$ input. Notably, our result shows that even a single Transformer can still achieve the same precision complexity as the class of all Transformers does. In particular, $\varGamma$ can decide any $\mathsf{P}$ language within log-precision.
@@ 1.1 Related work
Existing theoretical studies on Transformers fall under the classic one-model-one-task paradigm: they need to construct different Transformers for different tasks. There are two lines of related work: (i) when at most $\mathrm{O}(1)$ CoT steps are allowed, it has been shown that Transformers are capable but far from Turing-complete (Hahn, 2020; Hao et al., 2022; Liu et al., 2023a; Chiang et al., 2023; Merrill & Sabharwal, 2023; 2024b); (ii) when more CoT steps are allowed, it has been shown that the expressive power of Transformers increases with the number of CoT steps (Pérez et al., 2019; Bhattamishra et al., 2020; Pérez et al., 2021; Roberts, 2023; Merrill & Sabharwal, 2024a; Hou et al., 2024; Li et al., 2024). Besides that, there have recently been studies on the learnability (Malach, 2023; Grau-Moya et al., 2024) and the in-context learning capability (Akyürek et al., 2022; von Oswald et al., 2023; Zhang et al., 2024; Ahn et al., 2024; Vladymyrov et al., 2024). Nevertheless, no existing work studies the LLM prompting paradigm (i.e., the one-model-many-tasks paradigm). Our work is the first to bridge this gap to the best of our knowledge.
@@ 1.2 Technical overview
A core step of our proof is to construct a new model of computation (called $2$-PTMs) that can be easily encoded into a prompt using a finite alphabet. Furthermore, we show that $2$-PTMs are not only Turing-complete but also nearly as efficient as Turing machines.
@@ 2.1 Theory of computation
Turing machines.
Turing machines (TMs; Turing, 1937a) are an abstract model of computation that defines the notion of computability (Church, 1936; Kleene, 1936; Turing, 1937b). An $m$-tape TM is defined by a septuple $(Q,q_{\text{start}},q_{\text{halt}},\varSigma,\textvisiblespace,m,\delta)$ where $Q$ is the finite set of states, $q_{\text{start}}\in Q$ is the initial state, $q_{\text{halt}}$ is the halting state, $\varSigma$ is the finite alphabet of the tapes, $\textvisiblespace\in\varSigma$ is the blank symbol, $m\in\mathbb{N}$ is the number of tapes, and $\delta:(Q\setminus\{q_{\text{halt}}\})\times\varSigma^{m}\rightharpoonup Q\times(\varSigma\times\{\mathtt{L},\mathtt{S},\mathtt{R}\})^{m}$ is the transition function. See, e.g., Arora & Barak (2009) for details. By the Church–Turing thesis (Church, 1936; Kleene, 1936; Turing, 1937b), a function $\varphi:\operatorname{dom}\varphi\to\{\mathtt{0},\mathtt{1}\}^{*}$ is said to be computable if there exists a TM that computes $\varphi(\boldsymbol{x})$ for all inputs $\boldsymbol{x}\in\operatorname{dom}\varphi$.
@@
Time complexity.
Let $n$ denote the length of the input. A function $t(n):\mathbb{N}\to\mathbb{R}_{+}$ is said to be a complexity function iff $t(n)$ is non-decreasing and either $t(n)$ is a constant $>1$ or $t(n)\to\infty$ as $n\to\infty$. In this work, we refer to the time complexity as the the time complexity on a random-access machine (RAM) unless otherwise stated (e.g., when working with TMs). If $\Theta(t(n))$ is a complexity function, let $\mathsf{TIME}_{m}(t(n))$ denote the class of functions that can be computed by an $m$-tape TM within $\mathrm{O}(t(n))$ steps, let $\mathsf{TIME}(t(n)):=\bigcup{}_{m\geq 1}\mathsf{TIME}_{m}(t(n))$ denote the class of functions that can be computed by a TM within $\mathrm{O}(t(n))$ steps, and let $\mathsf{P}\subset\mathsf{TIME}(\operatorname{poly}(n))$ denote the class of (indicator functions of) languages that can be decided in polynomial time.
@@ 2.2 Neural networks
ReLU neural networks.
A ReLU neural network $\psi:\mathbb{R}^{e_{0}}\to\mathbb{R}^{e_{L}}$ is a composition $\psi:=\psi_{L-1}\circ\cdots\circ\psi_{0}$ of basic operations $\psi_{l}:\mathbb{R}^{e_{l}}\to\mathbb{R}^{e_{l+1}}$ ($l=0,\dots,L-1$) where each basic operation $\psi_{l}(\boldsymbol{z})$ ($\boldsymbol{z}\in\mathbb{R}^{e_{l}}$) is either an affine map $\psi_{l}(\boldsymbol{z}):=\boldsymbol{W}_{l}\boldsymbol{z}+\boldsymbol{b}_{l}$ ($\boldsymbol{W}_{l}\in\mathbb{R}^{e_{l+1}\times e_{l}}$, $\boldsymbol{b}_{l}\in\mathbb{R}^{e_{l+1}}$), an entry-wise ReLU activation $\psi_{l}(\boldsymbol{z}):=\max\{\boldsymbol{z},\boldsymbol{0}\}$ (Fukushima, 1969) with $e_{l+1}=e_{l}$, or a layer normalization $\psi_{l}(\boldsymbol{z}):=\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}1_{[\boldsymbol{z}\neq\boldsymbol{0}]}$ (Ba et al., 2016) with $e_{l+1}=e_{l}$.
Decoder-only Transformers.
Mainstream LLMs are based on decoder-only Transformers (Radford et al., 2018). Given a finite token alphabet $\varSigma$, a decoder-only Transformer $\varGamma:\varSigma^{+}\to\varSigma$ (with greedy decoding) is a composition $\varGamma:=\operatorname*{\arg\max}\circ\varGamma_{\text{out}}\circ\varGamma_{L-1}\circ\cdots\circ\varGamma_{0}\circ\varGamma_{\text{emb}}$ consists of an embedding layer $\varGamma_{\text{emb}}:\varSigma\to\mathbb{R}^{d}$, causal-attention Transformer layer $\varGamma_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ ($l=0,\dots,L-1$), and an output layer $\varGamma_{\text{out}}:\mathbb{R}^{d}\to\mathbb{R}^{\varSigma}$. A token embedding is a map $\operatorname{emb}:\varSigma\to\mathbb{R}^{d}$, and a positional encoding is a map $\operatorname{pos}:\mathbb{N}\to\mathbb{R}^{d}$. Following the common practice (Vaswani et al., 2017), given the input token sequence $\boldsymbol{v}=v_{0}\cdots v_{|\boldsymbol{v}|-1}\in\varSigma^{+}$, the embedding layer $\varGamma_{\text{emb}}$ adds a positional encoding $\operatorname{pos}(i)$ to each token embedding $\operatorname{emb}(v_{i})$:
@@
where $\operatorname{pos}$ is a computable function. Following existing theoretical works (e.g., Pérez et al., 2019; Hao et al., 2022; Merrill & Sabharwal, 2024a), we use $\operatorname{hardmax}$ attention as a realistic abstraction of $\operatorname{softmax}$. Given an $\mathbb{R}$ sequence $\boldsymbol{s}$, if the maximum value of $\boldsymbol{s}$ appears $t$ times, then we define $\operatorname{hardmax}_{i}(\boldsymbol{s}):=\frac{1}{t}$ if $s_{i}$ equals the maximum value and $\operatorname{hardmax}_{i}(\boldsymbol{s}):=0$ otherwise. Each Transformer layer $\varGamma_{l}$ has $H_{l}$ attention heads followed by a ReLU neural network $\phi_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Each attention head $k$ has a query map $\operatorname{qry}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a key map $\operatorname{qry}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a value map $\operatorname{val}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and a similarity map $\operatorname{sim}_{l,k}:\mathbb{R}\to\mathbb{R}$ ($k=0,\dots,H_{l}-1$), all of which are ReLU neural networks. The similarity scores between two tokens $v_{i}$ and $v_{j}$ are
@@
Algorithm 1 Autoregressive generation: $\operatorname{generate}_{\varGamma}(\boldsymbol{x})$
1:decoder-only Transformer $\varGamma:\varSigma^{+}\to\varSigma$; nonempty input $\boldsymbol{x}\in\varSigma^{+}$; stop token #$\texttt{\#}\in\varSigma$
2:generated output $\in\varSigma^{+}\cup\varSigma^{\upomega}$
3:repeat
4:     generate next token $c\leftarrow\varGamma(\boldsymbol{x})$
5:     append next token $\boldsymbol{x}\leftarrow\boldsymbol{x}\cdot c$
6:until $c=\texttt{\#}$
7:return generated output $\boldsymbol{x}$
@@ 3 Turing completeness of prompting
There exist a finite alphabet $\varSigma$, a finite-size decoder-only Transformer $\varGamma:\varSigma^{+}\to\varSigma$, and coding schemes $\operatorname{tokenize}:\{\mathtt{0},\mathtt{1}\}^{*}\to\varSigma^{*}$ and $\operatorname{readout}:\varSigma^{*}\to\{\mathtt{0},\mathtt{1}\}^{*}$ with which prompting is Turing-complete: for every computable function $\varphi:\operatorname{dom}\varphi\to\{\mathtt{0},\mathtt{1}\}^{*}$ with $\operatorname{dom}\varphi\subseteq\{\mathtt{0},\mathtt{1}\}^{*}$, there exists a prompt $\boldsymbol{\pi}_{\varphi}\in\varSigma^{+}$ such that for every input $\boldsymbol{x}\in\operatorname{dom}\varphi$, $\operatorname{generate}_{\varGamma}(\boldsymbol{\pi}_{\varphi}\cdot\operatorname{tokenize}(\boldsymbol{x}))$ computes a finite CoT, and
@@
Here, $\varSigma,\varGamma,\operatorname{tokenize},\operatorname{readout}$ are independent of $\varphi$; $\boldsymbol{\pi}_{\varphi}$ is independent of $\boldsymbol{x}$; and for any input $\boldsymbol{x}\in\{\mathtt{0},\mathtt{1}\}^{*}$, $\operatorname{tokenize}$ runs in $\mathrm{O}(|\boldsymbol{x}|)$ time, and $\operatorname{readout}$ runs in $\mathrm{O}(|\varphi(\boldsymbol{x})|)$ time.
@@ 3.2 Recording execution in CoT steps
It is known that finite-size Transformers without CoT steps are not universal (Hahn, 2020). To achieve Turing completeness, here we make use of CoT steps to record execution steps of the $2$-PTM $\boldsymbol{\iota}$ so that the Transformer can restore the state of the $2$-PTM at any execution step. In this subsection, we focus on the case where the input is empty; we will describe how to incorporate the input in Section 3.3.
@@
Algorithm 2 Extracting output from CoT ($\operatorname{readout}$)
1:generated CoT $\boldsymbol{v}\in\varSigma^{+}$
2:extracted output $\boldsymbol{y}\in\{\mathtt{0},\mathtt{1}\}^{*}$
3:initialize the left index $i\leftarrow-2$
4:while $v_{i}\neq\texttt{:}$ do
5:     decrement the left index $i\leftarrow i-1$
6:end while
7:return extracted output $\boldsymbol{v}_{i+1:-1}$

### v2

# Ask, and it shall be given: On the Turing completeness of prompting

@@ 1 Introduction
Nevertheless, existing theoretical studies fail to align with the LLM prompting practice (i.e., the one-model-many-tasks paradigm). In the era of LLMs, people train a single general-purpose LLM and provide the LLM with different prompts to perform different tasks. Since the success of GPT (Brown et al., 2020), the LLM prompting paradigm has revolutionized machine learning (Liu et al., 2023b). For example, a bonus capability arising from prompting is zero-shot learning (Wei et al., 2022a): when provided with suitable prompts, LLMs can even perform novel tasks that are not present in their training corpora. Such empirical success calls for a theoretical understanding of the LLM prompting paradigm:
@@
We answer this call and present the first theory on the LLM prompting paradigm to the best of our knowledge. In this work, we show that prompting is in fact Turing-complete: there exists a finite-size Transformer such that for any computable function, there exists a corresponding prompt following which the Transformer computes the function. Furthermore, we show that prompting is not only universal but also efficiently universal: even though we use one finite-size Transformer, it can still achieve nearly the same complexity bounds as that of the class of all unbounded-size Transformers.
@@
  • Simple construction. In fact, it is not hard to deduce the existence of such a Transformer $\varGamma$ by combining Theorem 1 of Hennie & Stearns (1966) and Theorem 3.4 of Pérez et al. (2019), but explicitly constructing the Transformer via that approach is cumbersome. Instead, this paper provides a simple constructive proof, which makes it easy to further study the properties of the construction such as CoT complexity and precision complexity.
  • CoT complexity. We show that our $\varGamma$ can compute any $\mathsf{TIME}_{2}(t(n))$ function within $\mathrm{O}(t(n))$ CoT steps and can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(t(n)\log t(n))$ CoT steps for any length-$n$ input. Notably, our result shows that even a single Transformer can still achieve nearly the same CoT complexity as the class of all Transformers does.
  • Precision complexity. We show that our $\varGamma$ can compute any $\mathsf{TIME}(t(n))$ function within $\mathrm{O}(\log(n+t(n)))$ bits of precision for any length-$n$ input. Notably, our result shows that even a single Transformer can still achieve the same precision complexity as the class of all Transformers does. In particular, $\varGamma$ can decide any $\mathsf{P}$ language within log-precision.
@@ 1.1 Related work
In modern machine learning (Wei et al., 2024; Chen et al., 2024; Liu et al., 2024a; b; c; 2023c; Qiu et al., 2024b; a; 2023; 2022; Xu et al., 2024; Qiu & Tong, 2024; Zeng et al., 2024; Lin et al., 2024; Yoo et al., 2025; 2024; Chan et al., 2024; Wu et al., 2024; He et al., 2024; Wang et al., 2023), Transformers have nowadays become a mainstream architecture. Existing theoretical studies on Transformers fall under the classic one-model-one-task paradigm: they need to construct different Transformers for different tasks. There are two lines of related work: (i) when at most $\mathrm{O}(1)$ CoT steps are allowed, it has been shown that Transformers are capable but far from Turing-complete (Hahn, 2020; Hao et al., 2022; Liu et al., 2023a; Chiang et al., 2023; Merrill & Sabharwal, 2023; 2024b); (ii) when more CoT steps are allowed, it has been shown that the expressive power of Transformers increases with the number of CoT steps (Pérez et al., 2019; Bhattamishra et al., 2020; Pérez et al., 2021; Roberts, 2023; Merrill & Sabharwal, 2024a; Hou et al., 2024; Li et al., 2024). Besides that, there have recently been studies on the learnability (Malach, 2023; Grau-Moya et al., 2024) and the in-context learning capability (Akyürek et al., 2022; von Oswald et al., 2023; Zhang et al., 2024; Ahn et al., 2024; Vladymyrov et al., 2024). Nevertheless, no existing work studies the LLM prompting paradigm (i.e., the one-model-many-tasks paradigm). Our work is the first to bridge this gap to the best of our knowledge.
@@ 1.2 Technical overview
A core step of our constructive proof is to construct a new model of computation (called $2$-PTMs) that can be easily encoded into a prompt using a finite alphabet. Furthermore, we show that $2$-PTMs are not only Turing-complete but also nearly as efficient as Turing machines.
@@ 2.1 Theory of computation
Turing machines. Turing machines (TMs; Turing, 1937a) are an abstract model of computation that defines the notion of computability (Church, 1936; Kleene, 1936; Turing, 1937b). An $m$-tape TM is defined by a septuple $(Q,q_{\text{start}},q_{\text{halt}},\varSigma,\textvisiblespace,m,\delta)$ where $Q$ is the finite set of states, $q_{\text{start}}\in Q$ is the initial state, $q_{\text{halt}}$ is the halting state, $\varSigma$ is the finite alphabet of the tapes, $\textvisiblespace\in\varSigma$ is the blank symbol, $m\in\mathbb{N}$ is the number of tapes, and $\delta:(Q\setminus\{q_{\text{halt}}\})\times\varSigma^{m}\rightharpoonup Q\times(\varSigma\times\{\mathtt{L},\mathtt{S},\mathtt{R}\})^{m}$ is the transition function ($\mathtt{L}$: left; $\mathtt{S}$: stay; $\mathtt{R}$: right). See, e.g., Arora & Barak (2009) for details. By the Church–Turing thesis (Church, 1936; Kleene, 1936; Turing, 1937b), a function $\varphi:\operatorname{dom}\varphi\to\{\mathtt{0},\mathtt{1}\}^{*}$ is said to be computable if there exists a TM that computes $\varphi(\boldsymbol{x})$ for all inputs $\boldsymbol{x}\in\operatorname{dom}\varphi$.
@@
Time complexity. Let $n$ denote the length of the input. A function $t(n):\mathbb{N}\to\mathbb{R}_{+}$ is said to be a complexity function iff $t(n)$ is non-decreasing and either $t(n)$ is a constant $>1$ or $t(n)\to\infty$ as $n\to\infty$. In this work, we refer to the time complexity as the the time complexity on a random-access machine (RAM) unless otherwise stated (e.g., when working with TMs). If $\Theta(t(n))$ is a complexity function, let $\mathsf{TIME}_{m}(t(n))$ denote the class of functions that can be computed by an $m$-tape TM within $\mathrm{O}(t(n))$ steps, let $\mathsf{TIME}(t(n)):=\bigcup{}_{m\geq 1}\mathsf{TIME}_{m}(t(n))$ denote the class of functions that can be computed by a TM within $\mathrm{O}(t(n))$ steps, and let $\mathsf{P}\subset\mathsf{TIME}(\operatorname{poly}(n))$ denote the class of (indicator functions of) languages that can be decided in polynomial time.
@@ 2.2 Neural networks
ReLU neural networks. A ReLU neural network $\psi:\mathbb{R}^{e_{0}}\to\mathbb{R}^{e_{L}}$ is a composition $\psi:=\psi_{L-1}\circ\cdots\circ\psi_{0}$ of basic operations $\psi_{l}:\mathbb{R}^{e_{l}}\to\mathbb{R}^{e_{l+1}}$ ($l=0,\dots,L-1$) where each basic operation $\psi_{l}(\boldsymbol{z})$ ($\boldsymbol{z}\in\mathbb{R}^{e_{l}}$) is either an affine map $\psi_{l}(\boldsymbol{z}):=\boldsymbol{W}_{l}\boldsymbol{z}+\boldsymbol{b}_{l}$ ($\boldsymbol{W}_{l}\in\mathbb{R}^{e_{l+1}\times e_{l}}$, $\boldsymbol{b}_{l}\in\mathbb{R}^{e_{l+1}}$), an entry-wise ReLU activation $\psi_{l}(\boldsymbol{z}):=\max\{\boldsymbol{z},\boldsymbol{0}\}$ (Fukushima, 1969) with $e_{l+1}=e_{l}$, or a layer normalization $\psi_{l}(\boldsymbol{z}):=\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}1_{[\boldsymbol{z}\neq\boldsymbol{0}]}$ (Ba et al., 2016) with $e_{l+1}=e_{l}$.
Decoder-only Transformers. Mainstream LLMs are based on decoder-only Transformers (Radford et al., 2018). Given a finite token alphabet $\varSigma$, a decoder-only Transformer $\varGamma:\varSigma^{+}\to\varSigma$ (with greedy decoding) is a composition $\varGamma:=\operatorname*{\arg\max}\circ\varGamma_{\text{out}}\circ\varGamma_{L-1}\circ\cdots\circ\varGamma_{0}\circ\varGamma_{\text{emb}}$ consists of an embedding layer $\varGamma_{\text{emb}}:\varSigma\to\mathbb{R}^{d}$, causal-attention Transformer layer $\varGamma_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ ($l=0,\dots,L-1$), and an output layer $\varGamma_{\text{out}}:\mathbb{R}^{d}\to\mathbb{R}^{\varSigma}$. A token embedding is a map $\operatorname{emb}:\varSigma\to\mathbb{R}^{d}$, and a positional encoding is a map $\operatorname{pos}:\mathbb{N}\to\mathbb{R}^{d}$. Following the common practice (Vaswani et al., 2017), given the input token sequence $\boldsymbol{v}=v_{0}\cdots v_{|\boldsymbol{v}|-1}\in\varSigma^{+}$, the embedding layer $\varGamma_{\text{emb}}$ adds a positional encoding $\operatorname{pos}(i)$ to each token embedding $\operatorname{emb}(v_{i})$:
@@
where $\operatorname{pos}$ is a computable function. Following existing theoretical works (e.g., Pérez et al., 2019; Hao et al., 2022; Merrill & Sabharwal, 2024a), we use $\operatorname{hardmax}$ attention as a realistic abstraction of $\operatorname{softmax}$. Given an $\mathbb{R}$ sequence $\boldsymbol{s}$, if the maximum value of $\boldsymbol{s}$ appears $t$ times, then we define $\operatorname{hardmax}_{i}(\boldsymbol{s}):=\frac{1}{t}$ if $s_{i}$ equals the maximum value and $\operatorname{hardmax}_{i}(\boldsymbol{s}):=0$ otherwise. Each Transformer layer $\varGamma_{l}$ has $H_{l}$ attention heads followed by a ReLU neural network $\phi_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Each attention head $k$ has a query map $\operatorname{qry}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a key map $\operatorname{key}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d_{l,k}}$, a value map $\operatorname{val}_{l,k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and a similarity map $\operatorname{sim}_{l,k}:\mathbb{R}\to\mathbb{R}$ ($k=0,\dots,H_{l}-1$), all of which are ReLU neural networks. The similarity scores between two tokens $v_{i}$ and $v_{j}$ are
@@ 3 Turing completeness of prompting
There exist a finite alphabet $\varSigma$, a finite-size decoder-only Transformer $\varGamma:\varSigma^{+}\to\varSigma$, and coding schemes $\operatorname{tokenize}:\{\mathtt{0},\mathtt{1}\}^{*}\to\varSigma^{*}$ and $\operatorname{readout}:\varSigma^{*}\to\{\mathtt{0},\mathtt{1}\}^{*}$ with which prompting is Turing-complete: for every computable function $\varphi:\operatorname{dom}\varphi\to\{\mathtt{0},\mathtt{1}\}^{*}$ with $\operatorname{dom}\varphi\subseteq\{\mathtt{0},\mathtt{1}\}^{*}$, there exists a prompt $\boldsymbol{\pi}_{\varphi}\in\varSigma^{+}$ such that for every input $\boldsymbol{x}\in\operatorname{dom}\varphi$, $\operatorname{generate}_{\varGamma}(\boldsymbol{\pi}_{\varphi}\cdot\operatorname{tokenize}(\boldsymbol{x}))$ computes [^1] Following prior work (e.g., Pérez et al., 2019), we assume that every arithmetic computation in $\varGamma$ is exact. a finite CoT, and
@@
Here, $\varSigma,\varGamma,\operatorname{tokenize},\operatorname{readout}$ are independent of $\varphi$; $\boldsymbol{\pi}_{\varphi}$ is independent of $\boldsymbol{x}$; for any input $\boldsymbol{x}\in\{\mathtt{0},\mathtt{1}\}^{*}$, $\operatorname{tokenize}$ and $\operatorname{readout}$ run in $\mathrm{O}(|\boldsymbol{x}|)$ and $\mathrm{O}(|\varphi(\boldsymbol{x})|)$ time on a random access machine (RAM), respectively.
@@ 3.2 Recording execution in CoT steps
It is known that finite-size Transformers without CoT steps are not universal (Hahn, 2020). To achieve Turing completeness, here we leverage CoT steps to record execution steps of the $2$-PTM $\boldsymbol{\iota}$ so that the Transformer can restore the state of $\boldsymbol{\iota}$ at any execution step. In this subsection, we focus on the case where the input is empty; we will describe how to incorporate the input in Section 3.3.